Because of a storm, ground controllers instruct the pilot of a plane flying at an altitude of 4 miles to make a turn and climb to an altitude of 4.2 miles. The model for the path of the plane during this maneuver is
r(t) = 10 cos 10π t i + 10 sin 10 π t j + ( 4+ 4t )k , 0≤ t≤ 1/20
where t is the time in hours and r is the position vector.
i. Determine the speed of the plane
ii. Calculate αT and αN
1
Expert's answer
2019-11-20T12:58:06-0500

Speed of the Plane


Given,

The model for the path of the plane during this maneuver is


"r(t) = 10 cos 10\pi \space \vec i + 10 sin 10\pi t \space \vec j + (4 + 4t ) \vec k , \space 0\le t\le\frac {1}{20}"

Here t is the time in hours and


r is the position vector


(a)


We can calculate the speed of the plane by the formula

"||r' (t)||"


So, Differentiate the position vector r(t) with respect to t



"v(t) = r' (t) = \frac {d}{dt} (10 cos 10\pi \space \vec i + 10 sin 10\pi t \space \vec j + (4 + 4t ) \vec k )"

"= - 100\pi sin 10\pi \space t \space \vec i \space + \space 100 \pi cos 10\pi t \space \vec j + 4 \vec k"



Now the Speed of the plane =



"||v(t) || = ||r' (t) || = || - 100 \pi sin 10\pi \space t \space \vec i \space + \space 100\pi cos 10\pi t \space \vec j + 4 \vec k \space ||"


"= \sqrt {(- 100 \pi sin10\pi t)^2 + (100\pi cos 10\pi t)^2 + 4^2 }""= \sqrt {( 10000 \pi^2 sin^2 10\pi t) + (10000\pi^2 cos^2 10\pi t) + 16 }"


"= \sqrt {10000 \pi^2 \space [( sin^2 (10\pi t ) + cos^2 (10\pi t)] + 16}"



"||v(t) || = \sqrt {10000\pi^2 (1) + 16} = \sqrt {10000\pi^2 + 16}"

"Now \space the \space Speed \space of \space the \space plane = = 4 \times \sqrt {625\pi^2 + 1} = 314.025 \space miles/hr"

(b).


Now, we need to calculate the "α_{T} \space and \space α_{N}"


Here, formula for Tangential acceleration is given by


"α_{T} = \frac {d}{dt} [||v||] = \frac {v.α}{||v||}"

We know,

"α(t) = v'(t) = \frac {d}{dt} (- 100 \pi sin 10\pi \space t \space \vec i \space + \space 100 \pi cos 10\pi t \space \vec j + 4 \vec k )"

"= -1000 \pi^2 \space cos 10\pi t \vec i - 1000 \pi^2 sin 10\pi t \space \vec j"



"v.α = (- 100\pi sin 10\pi \space t \space \vec i \space + \space 100 \pi cos 10\pi t \space \vec j + 4 \vec k ).
(-1000 \pi^2 \space cos 10\pi t \vec i - 1000 \pi^2 sin 10\pi t \space \vec j)"


"= 100000\pi^3 sin (10\pi t) \space cos (10\pi t) - 100000 \pi^3 sin (10\pi t) cos (10\pi t) + 0"


"v.\u03b1 = 0"


So, Tangential acceleration is

"α_{T} = \frac {v.α}{||v||} = \frac {0} { 314.025} = 0"

Here,  formula for Normal components of acceleration is given by


"α_{N} = \frac {||v \times α||}{||v||}"


"v \times α = \begin{vmatrix}
 i & j & k \\
 -100 \pi sin (10\pi t) & 100\pi cos (10\pi t) & 4 \\
 -1000\pi^2 cos (10\pi t) & -1000 \pi^2 sin (10\pi t) & 0
\end{vmatrix}"


"= i (0 + 4000 \pi^2 sin (10\pi t)) - j (0 + 4000\pi^2 cos (10\pi t) \\+ k ( 100000 \pi^3 sin^2 (10\pi t + 100000 \pi^3 cos^2 (10\pi t))"

.


"v \times α = = i ( 4000 \pi^2 sin (10\pi t)) - j ( 4000\pi^2 cos (10\pi t) \\+ k \times 100000 \pi^3 (sin^2 (10\pi t + cos^2 (10\pi t)))"


"v \times α = i ( 4000 \pi^2 sin (10\pi t)) - j ( 4000\pi^2 cos (10\pi t) \\+ k \times 100000 \pi^3"


"|| v \times α ||= \sqrt {(4000 \pi^2)^2 sin ^2(10\pi t) + (4000 \pi^2)^2 cos^2 (10\pi t) + 10000^2 \pi^6}"

"|| v \times α ||= \sqrt {(4000 \pi^2)^2 \space [sin ^2(10\pi t) + cos^2 (10\pi t)] + 10000^2 \pi^6}"

"|| v \times α ||= \sqrt {(4000 \pi^2)^2 \space + 10000^2 \pi^6}"



"|| v \times α ||= 4000 \pi^2 \sqrt {1 +625 \pi^2}"

"α_{N} = \frac {||v \times α||}{||v||} = \frac {4000\pi^2 \sqrt {1+ 625\pi^2}}{4 \times \sqrt {625\pi^2 + 1}}"

Normal components of acceleration is

"α_{N} = 1000 \pi^2"